1.(x-y+1)dx+(x+4y+1)dy=0
2.(6x-3y+2)dx+(2x-y-1)dy=0
1
Expert's answer
2020-12-29T17:15:21-0500

1) The given equation can be written as "\frac{dy}{dx}=-\frac{(x-y+1)}{(x+4y+1)}" ......(1)

Here the two lines "x-y+1=0" and "x+4y+1=0" intersect at the point "(-1,0)" .

We now transfer the origin by putting "x=-1+X," "y=0+Y" ......(2)

This gives "dx=dX" and "dy=dY"

Therefore from (1) , we get "\frac{dY}{dX}=-\frac{(X-Y)}{(X+4Y)}" ........(3)

Put "Y=vX." Then the equation (3) becomes

"v+X.\frac{dv}{dX}=-\frac{(1-v)}{(1+4v)}"

"\implies X.\frac{dv}{dX}=-\frac{(1-v)}{(1+4v)}-v"

"\implies X.\frac{dv}{dX}=-\frac{(1+4v^2)}{(1+4v)}"

Separating the variables, we get

"\frac{(1+4v)}{(1+4v^2)}dv+\frac{dX}{X}=0"

Now integrating we get,

"\intop\frac{(1+4v)}{(1+4v^2)}dv+\intop \frac{dX}{X}=0"

"\implies \intop\frac{1}{(1+4v^2)}dv+\intop\frac{4v}{(1+4v^2)}dv+\intop \frac{dX}{X}=0"

"\implies \frac {1}{2}.tan^{-1}(2v)+\frac{1}{2}.log|4v^2+1|+log|X|=log|c|" "["where "log|c|" is an integrating constant"]"

"\implies tan^{-1}(2v)+log|4v^2+1|+2log|X|=2log|c|"

"\implies \tan^{-1}(2.\frac{Y}{X})+log|\frac{4Y^2+X^2}{X^2}|+log|X|^2=log|c|^2" [Putting the value of "v=\frac{Y}{X}"]

"\implies \tan^{-1}(2.\frac{Y}{X})+log|{4Y^2+X^2}|-log|{X^2}|+log|X|^2=log|c|^2"

"\implies \tan^{-1}(2.\frac{Y}{X})+log|{4Y^2+X^2}|=K(say)"

"\implies \tan^{-1}(\frac{2y}{x+1})+log|{4y^2+(x+1)^2}|=K" [ Putting the value "X=x+1" and "Y=y" ]

Which is the required solution.


2) Here the two lines "6x-3y+2=0" and "2x-y-1=0" are parallel.

We see that the given equation may be written as "\frac{dy}{dx}=-[\frac{3(2x-y)+2}{(2x-y)-1}]." ........(1)

Put "2x-y=v" in (1) , so that "\frac{dv}{dx}=2-\frac{dy}{dx}"

Hence we obtain from (1), "2-\frac{dv}{dx}=-[\frac{3v+2}{v-1}]"

"\implies \frac{dv}{dx}=[\frac{3v+2}{v-1}]+2"

"\implies \frac{dv}{dx}=[\frac{5v}{v-1}]"

Separating the variables we get , "\frac{v-1}{5v}dv=dx" .......(2)

Now integrating (2) we have , "\intop \frac{v-1}{5v}dv=\intop dx"

"\implies \frac{1}{5}\intop dv- \frac{1}{5}\intop \frac{1}{v}dv=\intop dx"

"\implies \frac{1}{5}v-\frac{1}{5}log|v|=x+C" [where "C" is an integrating constant]

"\implies v-log|v|=5x+5C"

"\implies (2x-y)-log|2x-y|=5x+5C" [ Putting "v=2x-y" ]

"\implies log|2x-y|+(3x+y)=K" (say)

Which is the required solution.